Let Marginal utility of A (MUA)=2z-20-2x and Marginal utility of B (MUB)=2z=42-y. where z is marginal utility per Naira measured in utils, x is the amount spent on product A. and y is the amount spent on product B. Assume that the consumer has N10 to spend on A and B-that is, x+y= 10. i. How is the benefit of N10 maximized between A and B? ii. How much utility will the marginal Naira yield?
1
Expert's answer
2021-10-25T17:52:49-0400

Solution:

i.). The benefit of N10 between A and B is maximized where:"\frac{MU_{A} }{P_{A}} = \frac{MU_{B} }{P_{B}}"

= "\frac{2z - 20 - 2x}{x} = \frac{2z - 42 + y}{y}"


X = "\frac{2y(z-10)}{2z - 42+y}"


10 = x + y


"10 = \frac{2y(z-10)}{2z - 42+y} +y"

 

ii.). Marginal Naira will yield:


z = "\frac{420-92y+3y^{2} }{4(y-5)}"